# Minimally Entangled Typical Thermal States for Classical and Quantum Simulation of Gauge Theories at Finite Temperature and Density

### v1

@@ Abstract
Simulating strongly coupled gauge theories at finite temperature and density is a longstanding challenge in nuclear and high-energy physics that also has fundamental implications for condensed matter physics. In this work, we investigate the utility of minimally entangled typical thermal state (METTS) approaches to facilitate both classical and quantum computational studies of such systems. METTS techniques combine classical random sampling with imaginary time evolution, which can be performed on either a classical or a quantum computer, to estimate thermal averages of observables. We study the simplest model of a confining gauge theory, namely $\mathbb{Z}_{2}$ gauge theory coupled to spinless fermionic matter in 1+1 dimensions, which can be directly mapped to a local quantum spin chain with two- and three-body interactions. We benchmark both a classical matrix-product-state implementation of METTS and a recently proposed adaptive variational approach to METTS that is a promising candidate for implementation on near-term quantum devices, focusing on the equation of state as well as on various measures of fermion confinement. Of particular importance is the choice of basis for obtaining new METTS samples, which impacts both the classical sampling complexity (a key factor in both classical and quantum simulation applications) and complexity of circuits used in the quantum computing approach. Our work sets the stage for future studies of strongly coupled gauge theories with both classical and quantum hardware.
@@ I Introduction
In this paper we consider such a model, namely a $\mathbb{Z}_{2}$ lattice gauge theory coupled to spinless fermions in 1+1 dimensions [28, 29, 27, 30, 31, 32]. Like QCD, this model exhibits chiral symmetry breaking, confinement and string tension [28, 31, 33]; it is also relevant for studies of nonequilibrium condensed-matter phenomena like quantum many-body scars [29, 34, 35, 36, 32] and Hilbert-space fragmentation [27]. We explore the finite-temperature and -density properties of this model, focusing on measures of confinement and on the equation of state relating internal energy and fermion density. Our study adopts the minimally entangled typical thermal states (METTS) approach [37, 38, 39, 40], which combines imaginary time evolution (ITE) with a statistical sampling procedure to estimate quantum statistical-mechanics averages. Although originally developed as a tensor-network method, METTS can also be recast as a quantum algorithm (QMETTS), in which a quantum computer is used to perform the ITE subroutine [41]. While many approaches to ITE on quantum computers are possible, our study focuses on a recently proposed adaptive [42] variational [43] approach to QMETTS (AVQMETTS) [44]. We perform systematic benchmarks of grand-canonical-ensemble calculations within METTS with a view towards both classical and quantum computing approaches, focusing on a matrix product state (MPS) approach for the former and on exact statevector calculations for the latter. The classical and quantum approaches have common systematics in that they rely on the same classical sampling procedure, where the choice of sampling basis has an impact on convergence. AVQMETTS additionally depends on a choice of the operator pool used to construct the variational ansatz state, which impacts both the accuracy and quantum resource cost of the simulation. In carefully examining these systematics, our study lays groundwork for future progress in both classical and quantum simulation of gauge theories.
@@
The remainder of the paper is organized as follows. In Sec. II, we define the $\mathbb{Z}_{2}$ gauge theory model and review how to perform grand-canonical-ensemble calculations within METTS. In Sec. III, we benchmark classical METTS calculations of the internal energy density $\epsilon$ and fermion density $n$, focusing in particular on how to choose the optimal sampling basis. We then present METTS calculations of the $\epsilon$-$n$ equation of state, as well as two measures of confinement: Friedel oscillations of the fermion density [28] and string-antistring distribution functions [33]. In Sec. IV, we review the AVQMETTS method before benchmarking its performance with respect to sampling basis and operator pool. We move on to show results for the equation of state and Friedel oscillations, which we argue are a robust probe of confinement for the relatively small system sizes accessible on today’s quantum computers. We end in Sec. V with an outlook for future work.
@@ II $\mathbb{Z}_{2}$ Lattice Gauge Theory and METTS
In practice, the METTS sampling is performed in parallel with $S_{\textrm{w}}$ independent random walks, each of which generates $S_{0}$ METTSs. This amounts to an ensemble of size $S=S_{\textrm{w}}S_{0}$. To remove memory of the initial conditions, the first few thermal steps (typically the first ten) are excluded from the statistical analysis. Furthermore, the choice of measurement basis for METTS collapse is flexible, and previous investigations have demonstrated that alternating between $x$- and $z$-basis measurements (abbreviated as $xz$-basis) between adjacent thermal steps drastically reduces autocorrelations of the random walk [38, 44]. Here we perform a more systematic study of the choice of collapse basis and adopt the best strategy for our specific applications.
@@ III.2 Equation of state
To study the equation of state for the model (2)—i.e., the functional relationship between energy density $\epsilon$ and particle density $n$—we perform METTS calculations of a $L=60$ model with chemical potential varying from $\mu=-1.0$ to $\mu=1.0$ with a step $0.025$. The results for the free-fermion limit $h=0$ are plotted in Fig. 2(a) at three inverse temperatures $\beta=5,10$ and $20$. Without the confining field, the energy-particle density curve is symmetric around half-filling, indicative of particle-hole symmetry. Generally, the thermal energy density increases with temperature, as the number of excited states within the accessible energy window $\sim 1/\beta$ increases. As a result, $\epsilon$ grows the most at half-filling, and becomes trivially temperature-independent at zero or full filling, where the relevant subspace dimension reduces to 2 and the ground state is two-fold degenerate. For reference, the analytical calculations for the same free-fermion model are also shown in Fig. 2(a) with dotted lines, which agree perfectly with the METTS results and confirm the convergence of METTS calculations. In Fig. 2(b), we show results for the equation of state with $h=0.1$. The finite confining field clearly breaks the particle-hole symmetry. Compared with the free-fermion results in Fig. 2(a), the energy density becomes significantly lower for $n<0.5$ and reaches maximal energy reduction at $n=0$. In contrast, the variation of $\epsilon$ for $n>0.5$ is much smaller, and remains the same at $n=1$. This can be understood by considering the $\mathbb{Z}_{2}$ symmetry breaking by the confining field, which promotes the configurations with large negative magnetizations (or equivalently, those with long anti-strings as defined in Sec. III.3) and further lowers the energy. Since the total magnetization magnitude (length of anti-string) in each particle number sector is bounded by $L-2\lfloor N/2\rfloor$, the impact of the confining field on the thermal statistics grows from large filling to small filling of the model, consistent with the curves shown in Fig. 2(b). Note that the standard errors of $\epsilon$ and $n$ for these METTS calculations are below $10^{-3}$, i.e. they are are smaller than the line width of the curves.
@@ III.3 Probes of confinement: Friedel oscillations and string length distributions
Our results on Friedel oscillations and (anti-)string statistics at finite temperature are consistent with those obtained in Ref. [33] with a complementary MPS approach based on purifications [48, 49, 50]. The purification method computes thermal averages by preparing a state in which each physical qubit is maximally entangled with an auxiliary qubit, then performing ITE on the physical qubits for an imaginary time $\beta/2$. This method avoids the sampling entailed in METTS at the expense of doubling the number of qubits, which limits the system sizes that are easily accessible. The METTS approach adopted here allows us to simulate chains of up to $L=60$ fermion sites as in Figs. 2 and 3, whereas the finite-temperature calculations in Ref. [33] were limited to systems of $\sim 36$ sites. This increase in the accessible system sizes may be useful when studying, e.g., finite-temperature confinement phase diagrams in higher dimensions, where reducing the qubit overhead is particularly important.
@@ IV AVQMETTS simulations
In this section, we turn to a quantum version of METTS, specifically the AVQMETTS approach [44], for finite-temperature simulations of the model (2). Here we focus on benchmarking AVQMETTS calculations of the 1D model, motivated by the intriguing physics and available exact solutions. This section is divided into three parts. We first briefly outline AVQMETTS method and its technical details. Secondly, we present our AVQMETTS numerical results concerning the finite-temperature behavior of the energy and particle densities. Finally, we give a numerical estimation of the system-size scaling for the AVQMETTS circuit complexity characterized by the number of CNOT gates, which is a useful resource estimate for near-term applications. Even though AVQMETTS is a quantum algorithm, all the following calculations are performed using a state-vector simulator, which provide important benchmarks for future AVQMETTS calculations on quantum processing units.
@@ IV.2 Benchmarking the impact of CPS collapse basis
Figure 6: Minimal autocorrelation between thermal steps and accuracy of thermal energy and particle density from AVQMETTS calculations on the model (2) at $L=12$ without a confining field ($h=0$), and with a chemical potential set to $\mu=0.55$, which gives 1/3 particle filling. (a) Thermal energy density $\epsilon$ as a function of thermal steps using the alternating $yz$-basis (blue circles) at $\beta=10$. The results using $xz$-basis (orange squares) are also shown for comparison, together with the exact thermal energy density (dashed line). Each thermal step result represents an average over $S_{w}=288$ independent random walks. (b) Same as (a), but for particle density $n$. (c) Thermal energy density as a function of $\beta$ from AVQMETTS calculations using $yz$-basis, in comparison with exact results. The AVQMETTS result at each temperature is an average over $S_{w}\times S_{0}=288\times 4$ samples. (d) Same as (c) but for particle density.
@@
Table 1 shows the spread of the energies $\langle H\rangle_{i}$ and particle numbers $\langle N\rangle_{i}$ characterized by $\mathcal{D}_{E}$ and $\mathcal{D}_{N}$ from both AVQITE (bold numbers) and exact-ITE calculations. We used the same $\mathcal{M}=288$ states as in Fig. 5 to generate this table. The values obtained from AVQITE calculations are very close to those from exact-ITE, which further corroborates the accuracy of AVQITE. Across two representative temperatures ($\beta=1,2$), one can see that calculations using the $y$ basis consistently produce the smallest $\mathcal{D}_{E}$ and $\mathcal{D}_{N}$, while results using $z$($x$)-basis states are smaller for $\mathcal{D}_{E}$ ($\mathcal{D}_{N}$).
@@
Even though the AVQITE simulations for the CPSs in the $y$ basis yields the narrower energy and particle number distributions relative to the exact thermal averages (which implies better sampling efficiency for AVQMETTS), in practice, alternating the measurement basis for METTS collapse is found to reduce autocorrelation effects between thermal steps [38, 44]. Given the better AVQITE simulation accuracy for CPSs in the $z$-basis as demonstrated in Table 1 and the associated smaller $\mathcal{D}_{E}$, we choose the $yz$-basis for AVQMETTS, where calculations alternate between $y$- and $z$-basis measurements at odd and even thermal steps, respectively. To demonstrate the effectiveness of the $yz$-basis for AVQMETTS simulations, we plot in Fig. 6(a,b) the energy and particle number densities as functions of the number of thermal steps at $\beta=10$. Minimal autocorrelation effects are observed, as the thermal averages already converge to the exact thermal expectation values (dashed lines) at the second thermal step subject to fluctuations tied to the sample size $S=288$. This suggests that it suffices to skip the first thermal step for each Markovian random walk in AVQMETTS to get a METTS ensemble to estimate thermal averages. We follow this procedure for the following calculations. For reference, we also show the results with the $xz$-basis as adopted in Refs. [38, 44], where the calculations use an alternating $x$- and $z$-basis for collapse. One can see a faster convergence for $yz$-basis calculations compared with $xz$ results, especially for the energy density. In Fig. 6(c,d) we show the AVQMETTS estimates of thermal energy and particle density as functions of $\beta$ with $S=288\times 4$ samples. They agree well with the exact results up to a relative error of at most $2.6\%\ (0.8\%)$ for energy (particle) density.
@@ IV.3 AVQMETTS simulations of energy and particle densities
We further present the spatial distribution of the particle density $\langle n_{i}\rangle$ of the model without confinement in Fig. 7(c), which shows Friedel oscillations with a number of peaks equal to the particle occupancy $\langle N\rangle$, as discussed in Sec. III.3. Here one can the AVQMETTS results (symbols) also agree well with the exact results (dashed lines), where the four peaks are well resolved. As above, the deviation becomes slightly bigger with increasing temperature due to our use of a fixed ensemble size. At $\beta=5$ where the oscillations are almost washed out, the maximal relative error is about $\Delta_{n_{i}}=4.2\%$. Here we set $\mu=-0.55$ such that the system has $4$ particles in the ground state, consistent with calculations in Sec. III.3. We also find a consistent accuracy for the particle density distribution with finite confining field ($h=0.1$), where the number of peaks reduces to two due to particle pairing, as shown in Fig. 7(d).
Figure 8: System-size dependence of the AVQMETTS ansatz circuit complexity as measured by the number of CNOT gates $N_{\textrm{CX}}$. (a) Growth of $N_{\textrm{CX}}$ as a function of the system size $L$ for the deconfined free-fermion model with $h=0$. The results are shown for the two alternating $y$- and $z$-measurement basis for METTS collapse at temperatures $\beta=1,2$. Here we select a chemical potential $\mu$ for each system size to maintain a total filling $\langle N\rangle=L/4$ in the ground state. (b) Same as (a) but with confining field $h=0.1$ and with the chemical potential chosen to maintain a constant ground-state particle number $\langle N\rangle=2$ for all $L$. Each data point corresponds to an average over 288 random CPSs in the $y$ or $z$-basis. Dashed lines are best fits of the data with the function $f(L)=aL^{b}$.

### v2

# Minimally Entangled Typical Thermal States for Classical and Quantum Simulation of 1+1-Dimensional $\mathbb{Z}_{2}$ Lattice Gauge Theory at Finite Temperature and Density

@@ Abstract
Simulating strongly coupled gauge theories at finite temperature and density is a longstanding challenge in nuclear and high-energy physics that also has fundamental implications for condensed matter physics. In this work, we use minimally entangled typical thermal state (METTS) approaches to facilitate both classical and quantum computational studies of such systems. METTS techniques combine classical random sampling with imaginary time evolution, which can be performed on either a classical or a quantum computer, to estimate thermal averages of observables. We study 1+1-dimensional $\mathbb{Z}_{2}$ gauge theory coupled to spinless fermionic matter, which maps onto a local quantum spin chain. We benchmark both a classical matrix-product-state implementation of METTS and a recently proposed adaptive variational approach that is a promising candidate for implementation on near-term quantum devices, focusing on the equation of state as well as on various measures of fermion confinement. Of particular importance is the choice of basis for obtaining new METTS samples, which impacts both the classical sampling complexity (a key factor in both classical and quantum simulation applications) and complexity of circuits used in the quantum computing approach. Our work sets the stage for future studies of strongly coupled gauge theories with both classical and quantum hardware.
@@ I Introduction
In this paper we consider such a model, namely a $\mathbb{Z}_{2}$ lattice gauge theory coupled to spinless fermions in 1+1 dimensions [28, 29, 27, 30, 31, 32]. Like QCD, this model exhibits chiral symmetry breaking, confinement and string tension [28, 31, 33]; it is also relevant for studies of nonequilibrium condensed-matter phenomena like quantum many-body scars [29, 34, 35, 36, 32] and Hilbert-space fragmentation [27]. We explore the finite-temperature and -density properties of this model, focusing on measures of confinement and on the equation of state relating internal energy and fermion density. Our study adopts the minimally entangled typical thermal states (METTS) approach [37, 38, 39, 40, 41, 42], which combines imaginary time evolution (ITE) with a statistical sampling procedure to estimate quantum statistical-mechanics averages. Although originally developed as a tensor-network method, METTS can also be recast as a quantum algorithm (QMETTS), in which a quantum computer is used to perform the ITE subroutine [43]. While many approaches to ITE on quantum computers are possible, our study focuses on a recently proposed adaptive [44] variational [45] approach to QMETTS (AVQMETTS) [46]. We perform systematic benchmarks of grand-canonical-ensemble calculations within METTS with a view towards both classical and quantum computing approaches, focusing on a matrix product state (MPS) approach for the former and on exact statevector calculations for the latter. The classical and quantum approaches have common systematics in that they rely on the same classical sampling procedure, where the choice of sampling basis has an impact on convergence. AVQMETTS additionally depends on a choice of the operator pool used to construct the variational ansatz state, which impacts both the accuracy and quantum resource cost of the simulation. In carefully examining these systematics, our study lays groundwork for future progress in both classical and quantum simulation of gauge theories.
@@
The remainder of the paper is organized as follows. In Sec. II, we define the $\mathbb{Z}_{2}$ gauge theory model and review how to perform grand-canonical-ensemble calculations within METTS. In Sec. III, we benchmark classical METTS calculations of the internal energy density $\epsilon$ and fermion density $n$, focusing in particular on how to choose the optimal sampling basis. We then present METTS calculations of the $\epsilon$-$n$ equation of state, as well as two measures of confinement: Friedel oscillations of the fermion density [28] and string-antistring distribution functions [33]. In Sec. IV, we review the AVQMETTS method before benchmarking its performance with respect to sampling basis and operator pool. We move on to show results for the equation of state and Friedel oscillations, which we argue are a robust probe of confinement for the relatively small system sizes accessible on today’s quantum computers. In addition, we perform statevector and tensor-network simulations of the AVQMETTS algorithm and study the system-size scaling of its circuit complexity up to a system size of $L=32$, observing a power-law dependence. We end in Sec. V with an outlook for future work.
@@ II $\mathbb{Z}_{2}$ Lattice Gauge Theory and METTS
In practice, the METTS sampling is performed in parallel with $S_{\textrm{w}}$ independent random walks, each of which generates $S_{0}$ METTSs. This amounts to an ensemble of size $S=S_{\textrm{w}}S_{0}$. To remove memory of the initial conditions, the first few thermal steps (typically the first ten) are excluded from the statistical analysis. Furthermore, the choice of measurement basis for METTS collapse is flexible, and previous investigations have demonstrated that alternating between $x$- and $z$-basis measurements (abbreviated as $xz$-basis) between adjacent thermal steps drastically reduces autocorrelations of the random walk [38, 40, 46]. Here we perform a more systematic study of the choice of collapse basis and adopt the best strategy for our specific applications.
@@ III.2 Equation of state
To study the equation of state for the model (2)—i.e., the functional relationship between energy density $\epsilon$ and particle density $n$—we perform METTS calculations of a $L=60$ model with chemical potential varying from $\mu=-1.0$ to $\mu=1.0$ with a step $0.025$. The results for the free-fermion limit $h=0$ are plotted in Fig. 2(a) at three inverse temperatures $\beta=5,10$ and $20$. Without the confining field, the energy-particle density curve is symmetric around half-filling, indicative of particle-hole symmetry. Generally, the thermal energy density increases with temperature, as the number of excited states within the accessible energy window $\sim 1/\beta$ increases. As a result, $\epsilon$ grows the most at half-filling, and becomes trivially temperature-independent at zero or full filling, where the relevant subspace dimension reduces to 2 and the ground state is two-fold degenerate. For reference, the analytical calculations for the same free-fermion model are also shown in Fig. 2(a) with dotted lines, which agree perfectly with the METTS results and confirm the convergence of METTS calculations. In Fig. 2(b), we show results for the equation of state with $h=0.1$. The finite confining field clearly breaks the particle-hole symmetry. Compared with the free-fermion results in Fig. 2(a), the energy density becomes significantly lower for $n<0.5$ and reaches maximal energy reduction at $n=0$. In contrast, the variation of $\epsilon$ for $n>0.5$ is much smaller, and remains the same at $n=1$. This can be understood by considering the $\mathbb{Z}_{2}$ symmetry breaking by the confining field, which promotes the configurations with large negative magnetizations (or equivalently, those with long anti-strings as defined in Sec. III.3) and further lowers the energy. Since the total magnetization magnitude (length of anti-string) in each particle number sector is bounded by $L-2\lfloor N/2\rfloor$, the impact of the confining field on the thermal statistics grows from large filling to small filling of the model, consistent with the curves shown in Fig. 2(b). Note that the standard errors of $\epsilon$ and $n$ for these METTS calculations are below $10^{-3}$, i.e. they are are smaller than the line width of the curves. Furthermore, we have checked by performing additional METTS calculations at $L=120$ that increasing the system size results in an almost imperceptible change of the equation of state curves shown in Fig. 2(b). This builds confidence that these results are already close to the thermodynamic limit for $L\lesssim 60$.
@@ III.3 Probes of confinement: Friedel oscillations and string length distributions
Our results on Friedel oscillations and (anti-)string statistics at finite temperature are consistent with those obtained in Ref. [33] with a complementary MPS approach based on purifications [50, 51, 52]. The purification method computes thermal averages by preparing a state in which each physical qubit is maximally entangled with an auxiliary qubit, then performing ITE on the physical qubits for an imaginary time $\beta/2$. This method avoids the sampling entailed in METTS at the expense of doubling the number of qubits, which limits the system sizes that are easily accessible. The METTS approach adopted here allows us to simulate chains of up to (and even beyond) $L=60$ fermion sites as in Figs. 2 and 3, whereas the finite-temperature calculations in Ref. [33] were limited to systems of $\sim 36$ sites. This increase in the accessible system sizes may be useful when studying, e.g., finite-temperature confinement phase diagrams in higher dimensions, where reducing the qubit overhead is particularly important.
@@ IV AVQMETTS simulations
In this section, we turn to a quantum version of METTS, specifically the AVQMETTS approach [46], for finite-temperature simulations of the model (2). Here we focus on benchmarking AVQMETTS calculations of the 1D model, motivated by the intriguing physics and available exact solutions. This section is divided into three parts. We first briefly outline AVQMETTS method and its technical details. Secondly, we present our AVQMETTS numerical results concerning the finite-temperature behavior of the energy and particle densities. Finally, we give a numerical estimation of the system-size scaling for the AVQMETTS circuit complexity characterized by the number of CNOT gates, which is a useful resource estimate for near-term applications. Even though AVQMETTS is a quantum algorithm, all the following calculations are performed using classical simulations, which provide important benchmarks for future AVQMETTS calculations on quantum processing units.
@@ IV.2 Benchmarking the impact of CPS collapse basis
Figure 6: Minimal autocorrelation between thermal steps and accuracy of thermal energy and particle density from AVQMETTS calculations on the model (2) at $L=12$ without a confining field ($h=0$), and with a chemical potential set to $\mu=0.55$, which gives 1/3 particle filling. (a) Thermal energy density $\epsilon$ as a function of thermal steps using the alternating $yz$-basis (blue circles) at $\beta=10$. The results using $xz$-basis (orange squares) are also shown for comparison, together with the exact thermal energy density (dashed line). Each thermal step result represents an average over $S_{w}=288$ independent random walks.w (b) Same as (a), but for particle density $n$. (c) Thermal energy density as a function of $\beta$ from AVQMETTS calculations using $yz$- and $xz$-bases, in comparison with exact results. The AVQMETTS result at each temperature is an average over $S_{w}\times S_{0}=288\times 4$ samples. (d) Same as (c) but for particle density.
@@
Table 1 shows the spread of the energies $\langle H\rangle_{i}$ and particle numbers $\langle N\rangle_{i}$ characterized by $\mathcal{D}_{E}$ and $\mathcal{D}_{N}$ from both AVQITE (bold numbers) and exact-ITE calculations. We used the same $\mathcal{M}=288$ states as in Fig. 5 to generate this table. The values obtained from AVQITE calculations are very close to those from exact ITE, which further corroborates the accuracy of AVQITE. Across two representative temperatures ($\beta=1,2$), one can see that calculations using the $y$ basis consistently produce the smallest $\mathcal{D}_{E}$ and $\mathcal{D}_{N}$, while results using $z$($x$)-basis states are smaller for $\mathcal{D}_{E}$ ($\mathcal{D}_{N}$).
@@
Even though the AVQITE simulations for the CPSs in the $y$ basis yields the narrower energy and particle number distributions relative to the exact thermal averages (which implies better sampling efficiency for AVQMETTS), in practice, alternating the measurement basis for METTS collapse is found to reduce autocorrelation effects between thermal steps [38, 46]. Given the better AVQITE simulation accuracy for CPSs in the $z$-basis as demonstrated in Table 1 and the associated smaller $\mathcal{D}_{E}$, we choose the $yz$-basis for AVQMETTS, where calculations alternate between $y$- and $z$-basis measurements at odd and even thermal steps, respectively. To demonstrate the effectiveness of the $yz$-basis for AVQMETTS simulations, we plot in Fig. 6(a,b) the energy and particle number densities as functions of the number of thermal steps at $\beta=10$. Minimal autocorrelation effects are observed, as the thermal averages already converge to the exact thermal expectation values (dashed lines) at the second thermal step subject to fluctuations tied to the sample size $S=288$. This suggests that it suffices to skip the first thermal step for each Markovian random walk in AVQMETTS to get a METTS ensemble to estimate thermal averages. We follow this procedure for the following calculations. For reference, we also show the results with the $xz$-basis as adopted in Refs. [38, 46], where the calculations use an alternating $x$- and $z$-basis for collapse. One can see a faster convergence for $yz$-basis calculations compared with $xz$ results, especially for the energy density. In Fig. 6(c,d) we show the AVQMETTS estimates of thermal energy and particle density as functions of $\beta$ with $S=288\times 4$ samples. The data from both collapse bases ($yz$ and $xz$) agree well with the exact results. The maximum relative error for the energy density is about $2.6\%$ for the $yz$-basis and $2.0\%$ for the $xz$-basis, whereas the relative error for the particle density is $0.8\%$ for the former and $2.5\%$ for the latter.
@@ IV.3 AVQMETTS simulations of energy and particle densities
We further present the spatial distribution of the particle density $\langle n_{i}\rangle$ of the model without confinement in Fig. 7(c), which shows Friedel oscillations with a number of peaks equal to the particle occupancy $\langle N\rangle$, as discussed in Sec. III.3. Here one can verify that the AVQMETTS results (symbols) also agree well with the exact results (dashed lines), where the four peaks are well resolved. As above, the deviation becomes slightly bigger with increasing temperature due to our use of a fixed ensemble size. At $\beta=5$ where the oscillations are almost washed out, the maximal relative error is about $\Delta_{n_{i}}=4.2\%$. Here we set $\mu=-0.55$ such that the system has $4$ particles in the ground state, consistent with calculations in Sec. III.3. We also find a consistent accuracy for the particle density distribution with finite confining field ($h=0.1$), where the number of peaks reduces to two due to particle pairing, as shown in Fig. 7(d).
Figure 8: System-size dependence of the AVQMETTS ansatz circuit complexity as measured by the number of CNOT gates $N_{\textrm{CX}}$. (a) Growth of $N_{\textrm{CX}}$ as a function of the system size $L$ for the deconfined free-fermion model with $h=0$. The results are shown for the two alternating $y$- and $z$-measurement basis for METTS collapse at temperatures $\beta=1,2$. Each data point corresponds to an average over 288 random CPSs obtained using statevector simulations, while the error bars indicate the standard deviation of the corresponding distribution. (b) Same as (a) but the averaging is performed over 30 random CPSs obtained using tensor network simulations. We checked that increasing the sample size to 60 random CPSs does not meaningfully alter the results. (c) Same as (b) but with the confining field $h=0.1$. In (a)-(c) we select a chemical potential $\mu$ for each system size to maintain a total filling $\langle N\rangle=L/4$ in the ground state. Dashed lines are best fits of the data with the function $f(L)=aL^{b}$.